# A Multi-scale Information Integration Framework for Infrared and Visible Image Fusion

### v1

@@ IV-C2 Evaluation of the adaptive weight
IV-C3 Quantitative ablation experiments
Table II reports the quantitative ablation experiments. The average values of ablation experiment results are consistent with the comparison experiment with state-of-the-art models, i.e. the average value of VIF, SCD, CC and $Q^{AB/F}$ of our MDA are highest compared with others, which indicates that our MDA preserves the best detail and edge information from the input image pairs.
V Conclusion
In this paper, we propose a multi-scale framework for infrared and visible image fusion. We construct the residual downsample block and residual reconstruction block for multi-scale feature extraction and image reconstruction. Dual attention fusion block is applied to aggregate the complementary information from infrared and visible branches based on the spatial and channel attention module. Lastly, the information retention degree is obtained by the statistic methods to control weights of the intensity and gradient items in the loss function, which are used to constrain the information transferred from the inputs. We validate the effectiveness of our method on the TNO and RoadScene datasets. Qualitative and quantitative results show that our method achieves comparable results compared with the state-of-the-art methods and best satisfies human visual perception.

### v2

@@ 4.5.2 Evaluation of the adaptive weight
4.5.3 Quantitative ablation experiments
Table 2 reports the quantitative ablation experiments. The average values of ablation experiment results are consistent with the comparison experiment with state-of-the-art models, i.e. the average value of VIF, SCD, CC, AG and $Q^{AB/F}$ of our MDA are highest compared with others, which indicates that our MDA preserves the best detail and edge information from the input image pairs. Without using the adaptive weight in the $L_{image}$ and $L_{patch}$, the value of AG has dropped, which indicates that the global contrast information obtains insufficient preservation, and so does the fusion via the summation and concatenation methods.
4.5.4 Evaluation of the $\alpha$, $\beta$, $\gamma$ and size of the $W$
We analyze the influence of hyperparameters $\alpha$, $\beta$, $\gamma$ and size of the sliding window $W$ in the loss function and the quantitative result on the TNO dataset is shown in Table 3. The $\alpha$, $\beta$ and $\gamma$ are the weight controlling $L_{feature}$, $L_{style}$ and $L_{patch}$ in the loss function separately. $W$ controlls the size of the sliding window of the $L_{patch}$. We choose fourteen sets of hyperparameter values to test the influence of the different hyperparameters in the loss function, which are $\alpha=1e-5$, $\alpha=1e-6$, $\alpha=1e-10$, $\alpha=1e-11$, $\beta=1e3$, $\beta=1e4$, $\beta=1e6$, $\beta=1e7$, $\gamma=5$, $\gamma=3$, $\gamma=1.5$, $\gamma=1$, $W=15$, $W=23$. The comparison result demonstrates that AG and $Q^{AB/F}$ exhibit relatively high sensitivity to hyperparameters, which correspond to the retention of edge information. The variation of metrics changes relatively small when the $\alpha$ and $\beta$ change, which means the fusion result is less influenced by the $L_{feature}$ and $L_{style}$ compared with $L_{image}$. On the contrary, the variation of metrics changes more significantly when the $\gamma$ changes, especially the value of the VIF, AG and $Q^{AB/F}$, which indicates the balance between the global and local complementary information measurement would change the preservation of the intensity and detail information.
Table 3: Quantitative results of the $\alpha$, $\beta$, $\gamma$ and size of the $W$ on the TNO dataset. The best performance are shown in bold, and the second and third best performance are shown in red and blue, respectively.
$\alpha=1e-5$ 6.847 0.776 1.652 0.045 2.950 0.510 0.253 9.285
$\alpha=1e-6$ 7.170 0.645 1.670 0.048 2.074 0.527 0.229 9.755
$\alpha=1e-10$ 7.302 0.833 1.825 0.049 2.974 0.539 0.287 10.040
$\alpha=1e-11$ 7.208 0.832 1.818 0.047 2.662 0.534 0.264 9.806
$\beta=1e7$ 7.245 0.759 1.766 0.049 2.349 0.521 0.259 9.935
$\beta=1e6$ 7.277 0.785 1.768 0.048 2.378 0.515 0.235 9.855
$\beta=1e4$ 7.282 0.789 1.765 0.048 2.379 0.513 0.234 9.825
$\beta=1e3$ 7.201 0.878 1.797 0.047 2.931 0.522 0.240 9.774
$\gamma=5$ 7.289 0.802 1.770 0.049 2.210 0.515 0.203 9.823
$\gamma=2.5$ 7.281 0.796 1.769 0.049 3.102 0.514 0.307 9.833
$\gamma=1.5$ 7.281 0.785 1.763 0.048 2.965 0.512 0.279 9.826
$\gamma=1$ 7.289 0.777 1.751 0.048 2.235 0.509 0.188 9.803
$W=23$ 7.291 0.787 1.760 0.048 2.164 0.512 0.197 9.819
$W=15$ 7.290 0.786 1.760 0.048 2.165 0.511 0.223 9.820
Ours 7.111 0.850 1.850 0.044 3.524 0.550 0.357 9.735
5 Conclusion
In this paper, we propose a multi-scale framework for infrared and visible image fusion. We construct the residual downsample block and residual reconstruction block for multi-scale feature extraction and image reconstruction. Dual attention fusion block is applied to aggregate the complementary information from infrared and visible branches based on the spatial and channel attention module. Lastly, the information retention degree is obtained by the statistic methods to control weights of the intensity and gradient items in the loss function, which are used to constrain the information transferred from the inputs. We validate the effectiveness of our method on the TNO, RoadScene and $M^{3}FD$ datasets. Qualitative and quantitative results show that our method achieves comparable results compared with the state-of-the-art methods and best satisfies human visual perception.
6 Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grants U21A20514.
7 Declaration of interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.